A red candle is 45 cm tall and burns 3 cm per hour, a white candle is 30 cm tall and burns 1-1/2 cm per hour, how tall will each candle be after 1 hour, 3 hours, 5 hours and 12 hours?
What is the formula for calculating each candle's height at each time and which candle lasts longer? Will the red and white candles ever be the same height at the same time if started at the same time?
Please explain.
THANK YOU SO MUCH!


Answer by ankor@dixie-net.com(12704)

A red candle is 45 cm tall and burns 3 cm per hour, a white candle is 30 cm tall and burns 1-1/2 cm per hour, how tall will each candle be after 1 hour, 3 hours, 5 hours and 12 hours?
:
What is the formula for calculating each candle's height at each time
Let t = time in hrs
:
Red: Candle ht = 45 - 3t
White: Candle ht = 30 - 1.5t
:
Substitute 1, 3, 5 and 12 in each equation
:
Example: using the Red candle burning 12 hrs
ht = 45 - 3(12)
ht = 45 - 36
ht = 9 cm
:
:
which candle lasts longer?
Red
45 - 3t = 0
-3t = -45
t = -45/-3
t = +15 hrs duration of the Red candle
:
White
30 - 1.5t = 0
-1.5t = -30
t = -30/-1.5
t = 20 hrs duration of the white candle
:
:
Will the red and white candles ever be the same height at the same time if started at the same time?
:
White ht = Red ht
30 - 1.5t = 45 - 3t
3t - 1.5t = 45 - 30
1.5t = 15
t = 15/1.5
t = 10 hrs they will be equal
:
You can determine that height that is, substitute 10 for t in either equation
and solve for ht